OpenStudy (anonymous):

Solve. x2 - x = 6. x = 3 and x = -2 x = -3 and x = 2 x = 6 and x = -1 x = -6 and x = 1 @Liltico

OpenStudy (anonymous):

hmm, x^2 -x =6 x^2 -x -6 = 0, x^2 - 3x + 2x - 6 =0 x (x-3) + 2 (x-3) = 0 (x+2)(x-3) = 0 so, x +2 = 0 => x = -2 and x -3 = 0 => x =3 First option is correct